### v1

@@ Abstract
Fejér monotonicity is a well-established property commonly observed in sequences generated by optimization algorithms. In this paper, we introduce an extension of this property, called Fejér* monotonicity, which was initially proposed in [SIAM J. Optim., 34(3), 2535–2556 (2024)]. We discuss and build upon the concept by exploring its behavior within Hilbert spaces, presenting an illustrative example and insightful results regarding weak and strong convergence. We also compare Fejér* monotonicity with other weak notions of Fejér-like monotonicity, to better establish the role of Fejér* monotonicity in optimization algorithms.
@@
Fejér monotonicity convex analysis optimization algorithms
MSC:
49M27 65K05 65B99 90C25
journal: JOTA
@@ 1 Introduction
A useful concept for analyzing the convergence of optimization algorithms is the notion of Fejér monotonicity. A sequence $(x^{k})_{k\in\mathds{N}}$ defined in a Hilbert space $\mathcal{H}$ is said to be Fejér monotone with respect to a nonempty set $M\subset\mathcal{H}$ if for all $x\in M$ and all $k\in\mathds{N}$ we have $\lVert x^{k+1}-x\rVert\leq\lVert x^{k}-x\rVert$. It is well-known that if $(x^{k})_{k\in\mathds{N}}$ is Fejér monotone with respect to $M$, then $(x^{k})_{k\in\mathds{N}}$ is bounded, and if all its weak cluster points belong to $M$, then the entire sequence $(x^{k})_{k\in\mathds{N}}$ converges weakly to some point lying in $M$; see, for instance, (Bauschke:1996, , Theorem 2.16), and Combettes:2009 for further details and historical notes.
The notion of Fejér monotonicity is named after Hungarian mathematician Lipót Fejér (1880-1959) who introduced this notion in Fejer:1922 and further developed in Fejer:1937 and later in Agmon:1954 ; Motzkin:1954 . Fejér was the force behind the highly successful Hungarian school of analysis. His influence on the development of mathematics in the last century should be acknowledged. Indeed, Fejér was the advisor of mathematicians such as John von Neumann, Paul Erdös, George Pólya, Pál Turán and Gábor Szegö.
Weaker notions of Fejér monotonicity have been considered in the literature. For instance, the concept of quasi-Fejér monotonicity, first presented in Ermolev:1969 and further elaborated in several works such as Iusem:1994 ; Combettes:2001a .
In this work, we explore a slightly different notion, called Fejér* monotonicity. It was first introduced inBehling:2024c as a key ingredient to show that the Circumcentered-Reflection Method (CRM) Behling:2024b , proposed by these authors, solves the Convex Feasibility Problem (CFP). The main difference with the usual Fejér monotonicity notion relies on the fact that now the nonincreasing distance property holds for the tail of the sequence, starting at some index which depends on the considered point $x\in M$.
This paper is organized as follows. We start in Section 2 with a brief discussion on the role of Fejér monotonicity in optimization algorithms. In Section 3 we present some basic properties of Fejér* monotonicity, and we make a discussion, upon an illustrative example on properties of Fejér* monotonicity. In Section 4 we compare Fejér* sequences and quasi-Fejér sequences. Finally, in Section 5 we present some concluding remarks.
@@ 2 Fejér* monotonicity in optimization algorithms
$\left\lVert x^{k+1}-x\right\rVert\leq\left\lVert x^{k}-x\right\rVert.$
We address now the role of Fejér and Fejér monotonicity in connection with fixed-point and feasibility problems. Consider an operator $T:\mathds{R}^{n}\to\mathds{R}^{n}$. We remind that $T$ is a contraction if there exists $\tau\in(0,1)$ such that
$\left\lVert T(x)-T(y)\right\rVert\leq\tau\left\lVert x-y\right\rVert$
for all $x,y\in\mathds{R}^{n}$.
It is well known, and elementary, that a contraction has a unique fixed point (i.e., a point $\bar{x}\in\mathds{R}^{n}$ such that $T(\bar{x})=\bar{x}$) and that the sequence obtained by iterating the operator (i.e., the sequence $(x^{k})_{k\in\mathds{N}}\subset\mathds{R}^{n}$ defined as $x^{k+1}=T(x^{k})$ starting at any $x^{0}\in\mathds{R}^{n}$), converges to such unique fixed point.
@@
$\left\lVert T(x)-T(y)\right\rVert\leq\left\lVert x-y\right\rVert$
for all $x,y\in\mathds{R}^{n}$.
The set of fixed points of a nonexpansive operator $T$ may be empty, or infinite, and the behavior of the sequence $(x^{k})_{k\in\mathds{N}}$ obtained by iterating $T$ may fail to converge, but given any fixed point $\bar{x}$ of $T$, it follows immediately from Section 2 that $\left\lVert x^{k+1}-\bar{x}\right\rVert\leq\left\lVert x^{k}-\bar{x}\right\rVert$, i.e., the sequence is Fejér monotone to the set $M$ of fixed points of $T$, and thus it enjoys the basic properties of Fejér monotone sequences, namely that it is bounded, and that if some of its cluster points belongs to $M$, then the sequence $(x^{k})_{k\in\mathds{N}}$ converges to it; see, for instance, Bauschke:2017a for more on nonexpansiveness and Fejér monotonicity.
Among the nonexpansive operators, projections onto convex sets are quite significant. We remind that, given a closed and convex set $C\subset\mathds{R}^{n}$, the projection onto $C$, $P_{C}:\mathds{R}^{n}\to C$, is defined as $P_{C}(x)\coloneqq\operatorname*{arg\,min}_{y\in C}\left\lVert x-y\right\rVert$. Convex combinations of projections (i.e., operators of the form $\bar{P}=\sum_{i+1}^{m}\alpha_{i}P_{C_{i}}$ with $\alpha_{i}\in[0,1],\sum_{i=1}^{m}\alpha_{i}=1$, and $C_{i}\subset\mathds{R}^{n}$ convex) are nonexpansive, and the same holds for compositions of projections, i.e., operator of the form $\widehat{P}=P_{C_{m}}\circ\dots\circ P_{C_{1}}$. The set of fixed points of a projection $P_{C}$ is $C$, and the sets of fixed points of $\bar{P}$ and $\widehat{P}$, as defined above are equal to the set $C:=\cap_{i=1}^{m}C_{i}$. Iterating the operators $\bar{P}$ and $\widehat{P}$, one obtains the so-called Simultaneous Projections and Sequential Projections algorithms, which aim to solve the (CFP), consisting of finding a point in the intersection $C$ of closed and convex sets $C_{1},\dots,C_{m}$. See Bauschke:1996 for a detailed discussion on these methods, which can be traced back to the seminal works of Cimmino Cimmino:1938 and Kaczmark Kaczmarz:1937 . In view of the discussion above, the sequences generated by these two methods are Fejér monotone to $C$.
@@
    $\left\lVert T(x)-T(y)\right\rVert^{2}\leq\left\lVert x-y\right\rVert^{2}-\left\lVert(T(x)-T(y))-(x-y)\right\rVert^{2}$
    for all $x,y\in\mathds{R}^{n}$; and
@@
    $\left\lVert T(x)-T(y)\right\rVert\leq\left\lVert x-y\right\rVert$
    for all $x,y\in\mathds{R}^{n}$, and $T(x)-T(y)=x-y$ whenever equality holds in (ii).
Projections onto convex sets, as well as the above defined operators $\bar{P}$ and $\widehat{P}$, are firmly nonexpansive. Convex combinations of firmly nonexpansive operators may fail to be firmly nonexpansive, but they are nonexpansive plus (see Arefidamghani:2023 ). With the help of the Fejér convergence properties and the notion of firm nonexpansiveness, it is easy to prove that the Simultaneous and the Successive Projection algorithms do converge to a point in $C$, whenever $C$ is nonempty (see Bauschke:1996 ).
@@
$C^{k}\subset C^{k+1}\subset\dots\subset C,$
for all $k$. For instance, if $C\coloneqq\{x\in\mathds{R}^{n}\mid g(x)\leq 0\}$ for some convex $g:\mathds{R}^{n}\to\mathds{R}$, then $C^{k}$ could be chosen as
$C^{k}\coloneqq\{x\in\mathds{R}^{n}\mid g(x)+\epsilon_{k}\leq 0\},$
@@
where $(\epsilon_{k})_{k\in\mathds{N}}$ is a sequence of positive real numbers decreasing to 00. Then one could consider the sequence $(x^{k})_{k\in\mathds{N}}$ given by $x^{k+1}=T_{k}(x^{k})$, where $T_{k}$ is the projection onto a convex set which contains $C^{k}$. In such a case, the nonexpansiveness of $T^{k}$ gives $\left\lVert x^{k+1}-y\right\rVert\leq\left\lVert x^{k}-y\right\rVert$ but only for points $y\in C^{k}$, not in the whole set $C$.
However, if we take $C^{k}$ as in Section 2, and $C$ has nonempty interior, any point $y\in$ int$(C)$ (i.e., such that $g(y)<0$), will belong to $C^{k}$ for all $k$ bigger than $k(y)$, defined as the first integer such that $\epsilon_{k(y)}<-g(y)$, due to Section 2, but this is exactly the same as saying that $(x^{k})_{k\in\mathds{N}}$ is Fejér monotone to int$(C)$.
Hence, Fejér* monotonicity is precisely what is needed for the analysis of algorithms which use approximation schemes as in Section 2. Among them, we mention the algorithms in Iusem:1986a ; Iusem:1986 for CFP. These methods employ the scheme in Section 2 combined with operators related to $\bar{P}$ and $\widehat{P}$, where the projections onto the sets $C_{i}$, approximated from the inside by sets $C^{k}_{i}$, as in Section 2, are replaced by projections onto hyperplanes separating $x^{k}$ from $C^{k}_{i}$. In these cases, the Fejér monotonicity ensures convergence to a point lying possibly in the boundary of $C$ (despite the fact that the sequence is Fejér monotone to int$(C)$). A similar situation occurs with the algorithm in Behling:2024c , where the algorithm in Iusem:1986a is enhanced through the addition of a circumcentering step, using the machinery developed during the last few years in Araujo:2022 ; Arefidamghani:2021 ; Arefidamghani:2023 ; Behling:2018 ; Behling:2018a ; Behling:2020 ; Behling:2021b ; Behling:2023 ; Behling:2024 ; Behling:2024b ; Behling:2024c .
@@
It is worthwhile to note that in the case of Iusem:1986 ; Iusem:1986a ; Behling:2024c , under a Slater hypothesis (i.e., when $\operatorname{int}(C)$ is nonempty), if the $\epsilon_{k}$’s decrease to 00 slowly enough, e.g., so that $\sum_{k=0}^{\infty}\epsilon_{k}=+\infty$, for instance $\epsilon_{k}=1/k$, convergence is indeed finite, i.e., there exists $\bar{k}$ such that $x^{\bar{k}}$ belongs to $C=\bigcap_{i=1}^{m}C_{i}$.
@@ 3 Properties of Fejér* monotonicity
The first result of this section was already proved in (Behling:2024c, , Prop. 2.5). However, for the sake of completeness, we present it here.
Proposition 1
Let $(x^{k})_{k\in\mathds{N}}\subset\mathcal{H}$ be a Fejér* monotone w.r.t. a nonempty set $M$ in $\mathcal{H}$. Then,
@@
    $(x^{k})_{k\in\mathds{N}}$ is Fejér* monotone w.r.t. $\operatorname{conv}(M)$.
Proof
@@
Item (ii) is a direct consequence of Definition 1, since the $N(x)$-tail of sequence $(\left\lVert x^{k}-x\right\rVert)_{k\in\mathds{N}}$ is monotone and bounded, so the sequence converges.
For item (iii), take any $w\in\operatorname{conv}(M)$. Thus, $w$ can be written as $w=\sum_{i=1}^{p}\lambda_{i}w^{i}$, where $w^{i}\in M$, $\lambda_{i}\in[0,1]$ for all $i=1,2,\ldots,p$, $\sum_{i=1}^{p}\lambda_{i}=1$ and $p\in\mathds{N}$. Taking into account that $(x^{k})_{k\in\mathds{N}}$ is Fejér* monotone w.r.t. $M$, for each $i=1,2,\ldots,p$, there exists $N(w^{i})$ such that $\left\lVert x^{k+1}-w^{i}\right\rVert\leq\left\lVert x^{k}-w^{i}\right\rVert$ for all $k\geq N(w^{i})$. Squaring both sides of the last inequality and rearranging the terms, we get
00 0 $\leq\left\lVert{x^{k}}\right\rVert^{2}-\left\lVert{x^{k+1}}\right\rVert^{2}-2\left\langle x^{k}-x^{k+1},w^{i}\right\rangle.$ (1)
The above inequality also holds for all $k\geq K\coloneqq\max\{N(w^{1}),N(w^{2}),\ldots,N(w^{p})\}$. Multiplying Eq. 1, for each $i=1,2,\ldots,p$, by the respective $\lambda_{i}$ and adding up, we obtain
00 0 $\leq\left\lVert{x^{k}}\right\rVert^{2}-\left\lVert{x^{k+1}}\right\rVert^{2}-2\left\langle x^{k}-x^{k+1},\sum_{i=1}^{p}\lambda_{i}w^{i}\right\rangle$ (2)
@@
and therefore, $\left\lVert x^{k+1}-w\right\rVert\leq\left\lVert x^{k}-w\right\rVert$, for all $k\geq K$. So $(x^{k})_{k\in\mathds{N}}$ is Fejér* monotone w.r.t. $\operatorname{conv}(M)$, as the proof is complete. ∎
We remark that a Fejér monotone sequence with respect to a nonempty set $M\subset\mathcal{H}$ is also Fejér monotone with respect to $\operatorname{cl}\left({\operatorname{conv}(M)}\right)$, the closure of $\operatorname{conv}(M)$; see (Bauschke:2023, , Lem. 2.1(ii)). Furthermore, Proposition 1(iii) guarantees that a Fejér* monotone sequence w.r.t. $M$ is as well Fejér* monotone w.r.t. $\operatorname{conv}(M)$. However, the Fejér* monotonicity cannot always be extended to $\operatorname{cl}\left({\operatorname{conv}(M)}\right)$, as shown in the next example.
Example 1
@@
$x^{2\ell+1}\coloneqq x^{2\ell}+(1/2^{\ell},0)\;,\quad x^{2\ell+2}\coloneqq\left(0,\sqrt{\|x^{2\ell+1}-(1,0)\|^{2}-1}\right).$
If $k$ is even, $x^{k+1}$ is obtained by moving $1/2^{(k/2)}$ horizontally to the right. If $k$ is odd, $x^{k+1}$ is the intersection between the vertical axis and the arc centered in $w^{0}=(1,0)$ passing through $x^{k}$; see Figure 1.
Figure 1: From Example 1: set $M$, first nine points of sequence $(x^{k})_{k\in\mathds{N}}$, $\operatorname{conv}(M)=(0,1]\times\{0\}$, $\operatorname{cl}\left({\operatorname{conv}(M)}\right)\setminus\operatorname{conv}(M)=\{(0,0)\}$ and limit point $x^{*}$ of sequence $(x^{k})_{k\in\mathds{N}}$.
We now show that $(x^{k})_{k\in\mathds{N}}$ is Fejér* with respect to $M$. Denote $x^{k}=(\alpha_{k},\beta_{k})$, $k\in\mathds{N}$. One can see that $1\leq\beta_{k+1}\leq\beta_{k}$ as well as $\|x^{k}-w^{0}\|\geq 1$, for all $k\in\mathds{N}$, so that the sequence $(x^{k})_{k\in\mathds{N}}$ is well-defined.

### v2

@@ Abstract
Fejér monotonicity is a well-established property often observed in sequences generated by optimization algorithms. In this paper, we study an extension of this property, called Fejér* monotonicity, which was initially proposed in [SIAM J. Optim., 34(3), 2535–2556 (2024)]. We discuss and explore its behavior within Hilbert spaces as a tool for optimization algorithms. Additionally, we investigate weak and strong convergence properties of this novel concept. Through illustrative examples and insightful results, we contrast Fejér* with weaker notions of quasi-Fejér-type monotonicity.
@@
Fejér* monotonicity, Fejér monotonicity, convex analysis, optimization algorithms
pacs:
MSC (2020)]49M27, 65K05, 90C25
@@ 1 Introduction
A useful concept for analyzing the convergence of optimization algorithms is the notion of Fejér monotonicity. We denote by $\mathds{N}$ the set of all nonnegative integers. A sequence $(x^{k})_{k\in\mathds{N}}$ defined in a real Hilbert space $\mathcal{H}$ is said to be Fejér monotone with respect to a nonempty set $M\subset\mathcal{H}$ if for all $x\in M$ and all $k\in\mathds{N}$ we have $\lVert x^{k+1}-x\rVert\leq\lVert x^{k}-x\rVert$. It is well-known that if $(x^{k})_{k\in\mathds{N}}$ is Fejér monotone with respect to $M$, then $(x^{k})_{k\in\mathds{N}}$ is bounded, and if all its weak cluster points belong to $M$, then the entire sequence $(x^{k})_{k\in\mathds{N}}$ converges weakly to some point lying in $M$; see, for instance, [Bauschke:1996, Thm 2.16], [Bauschke:2017a, Prop. 5.4 and Thm. 5.5, p. 92], and [Combettes:2009] for further details and historical notes.
The notion of Fejér monotonicity is named after Hungarian mathematician Lipót Fejér (1880-1959) who introduced this notion in [Fejer:1922] and further developed it in [Fejer:1937] and later in [Agmon:1954, Motzkin:1954]. Fejér was the force behind the highly successful Hungarian school of analysis. His influence on the development of mathematics in the last century should be acknowledged. Indeed, Fejér was the advisor of mathematicians such as John von Neumann, Paul Erdös, George Pólya, Pál Turán and Gábor Szegö.
Weaker notions of Fejér monotonicity have been considered in the literature. For instance, the concept of quasi-Fejér monotonicity, first presented in [Ermolev:1969] and further elaborated in several works such as [Iusem:1994, Combettes:2001a].
In this work, we explore a slightly different notion, called Fejér* monotonicity. It was first introduced in [Behling:2024c] as a key ingredient to show that the Circumcentered-Reflection Method (CRM) [Behling:2024b], proposed by the authors of [Behling:2024c], solves the Convex Feasibility Problem (CFP). The main difference with the usual Fejér monotonicity notion relies on the fact that now the nonincreasing distance property holds for the tail of the sequence, starting at some index which depends on the considered point $x\in M$.
In this work we further explore this novel concept. Indeed, the main contributions of this paper are twofold: (i) We present a detailed discussion on the properties of Fejér* monotonicity, including the ones that can be used as a tool to analyze convergence of sequences generated by optimization algorithms as well as the relationship of Fejér* with other weaker notions of Fejér monotonicity; (ii) We provide an illustrative example that highlights the differences between Fejér* and quasi-Fejér monotonicity. This example is key to clarify the connections and differences of these concepts.
This paper is organized as follows. We start in Section˜2 with a brief discussion on the role of Fejér monotonicity in optimization algorithms. In Section˜3 we present some basic properties of Fejér* monotonicity, and we discuss, upon illustrative examples, some properties of Fejér* monotonicity sequences. In Section˜4 we compare Fejér* sequences and quasi-Fejér sequences. Finally, in Section˜5 we present some concluding remarks.
@@ 2 Fejér* monotonicity in optimization algorithms
$\|x^{k+1}-x\|\leq\|x^{k}-x\|.$ (2)
Note that Fejér monotonicity implies Fejér* monotonicity, but the converse is not true in general. We address now the role of Fejér and Fejér* monotonicity in connection with fixed-point and feasibility problems. Consider an operator $T:\mathcal{H}\to\mathcal{H}$. We remind that $T$ is a contraction if there exists $\tau\in(0,1)$ such that
$\left\lVert T(x)-T(y)\right\rVert\leq\tau\left\lVert x-y\right\rVert,$ (3)
for all $x,y\in\mathcal{H}$. It is well known that a contraction $T$ has a unique fixed point (i.e., a point $\bar{x}\in\mathcal{H}$ such that $T(\bar{x})=\bar{x}$) and that the sequence obtained by iterating the operator $T$ (i.e., the sequence $(x^{k})_{k\in\mathds{N}}\subset\mathcal{H}$ defined as $x^{k+1}=T(x^{k})$ starting at any $x^{0}\in\mathcal{H}$), converges to such a unique fixed point; for further details on contraction mappings, see [Kirk:2001].
@@
$\left\lVert T(x)-T(y)\right\rVert\leq\left\lVert x-y\right\rVert,$ (4)
for all $x,y\in\mathcal{H}$.
The set of fixed points of a nonexpansive operator $T$ may be empty, or infinite, and the behavior of the sequence $(x^{k})_{k\in\mathds{N}}$ obtained by iterating $T$ may fail to converge, but given any fixed point $\bar{x}$ of $T$, it follows immediately from Eq.˜4 that $\left\lVert x^{k+1}-\bar{x}\right\rVert\leq\left\lVert x^{k}-\bar{x}\right\rVert$, i.e., the sequence is Fejér monotone with respect to the set $M$ of fixed points of $T$, and thus it enjoys the basic properties of Fejér monotone sequences, namely that it is bounded, and that if all its cluster points belong to $M$, then the sequence $(x^{k})_{k\in\mathds{N}}$ converges to one of them; see, for instance, [Bauschke:2017a] for more on nonexpansiveness and Fejér monotonicity.
Among the nonexpansive operators, projections onto convex sets are quite significant. We remind that, given a nonempty, convex and closed set $C\subset\mathcal{H}$, the projection onto $C$, $P_{C}:\mathcal{H}\to C$, is defined as $P_{C}(x)\coloneqq\operatorname*{arg\,min}_{y\in C}\left\lVert x-y\right\rVert$. Convex combinations of projections (i.e., operators of the form $\bar{P}=\sum_{i=1}^{m}\alpha_{i}P_{C_{i}}$ with $\alpha_{i}\in[0,1],\sum_{i=1}^{m}\alpha_{i}=1$, and $C_{i}\subset\mathcal{H}$ convex) are nonexpansive, and the same holds for compositions of projections, i.e., operator of the form $\widehat{P}=P_{C_{m}}\circ\dots\circ P_{C_{1}}$. The set of fixed points of a projection $P_{C}$ is $C$, and the sets of fixed points of $\bar{P}$ and $\widehat{P}$, as defined above, are equal to the set $C:=\cap_{i=1}^{m}C_{i}$ whenever $C\neq\emptyset$. Iterating the operators $\bar{P}$ and $\widehat{P}$, one obtains the so-called Simultaneous Projections and Sequential Projections algorithms, which aim to solve the CFP, consisting of finding a point in the intersection $C$ of nonempty, convex and closed sets $C_{1},\dots,C_{m}$. See the work of Bauschke:1996 for a detailed discussion on these methods, which can be traced back to the seminal works of Cimmino:1938 and Kaczmarz:1937. In view of the discussion above, the sequences generated by these two methods are Fejér monotone with respect to $C$.
@@
    $\left\lVert T(x)-T(y)\right\rVert^{2}\leq\left\lVert x-y\right\rVert^{2}-\left\lVert(T(x)-T(y))-(x-y)\right\rVert^{2}$ (5)
    for all $x,y\in\mathcal{H}$; and
@@
    $\left\lVert T(x)-T(y)\right\rVert\leq\left\lVert x-y\right\rVert$ (6)
    for all $x,y\in\mathcal{H}$, and T(x)T(y)=xyT(x)-T(y)=x-y whenever equality holds in Eq.˜6.
Projections onto nonempty, convex and closed sets, as well as the above defined operator $\bar{P}$, are firmly nonexpansive. With the help of Fejér convergence properties and the notion of firm nonexpansiveness, it is easy to prove that the Simultaneous and the Successive Projection algorithms do converge to a point in $C$, whenever $C$ is nonempty (see [Bauschke:1996]).
@@
$C^{k}\subset C^{k+1}\subset\dots\subset C,$ (7)
@@
for all $k$. Then one may consider the sequence $(x^{k})_{k\in\mathds{N}}$ given by $x^{k+1}\coloneqq T_{k}(x^{k})$, where $T_{k}$ is the projection onto a nonempty, convex and closed set which contains $C^{k}$ (but may not be contained or even contain $C$). In such a case, the nonexpansiveness of $T_{k}$ gives $\left\lVert x^{k+1}-y\right\rVert\leq\left\lVert x^{k}-y\right\rVert$, but only for points $y\in C^{k}$, not in the whole set $C$.
For instance, suppose that we have $C\coloneqq\{x\in\mathcal{H}\mid g(x)\leq 0\}$, for some function $g:\mathcal{H}\to\mathds{R}$ that is strictly convex. This choice is consistent with the Slater hypothesis (i.e., the existence of $\hat{y}\in\mathcal{H}$, such that $g(\hat{y})<0$) that is needed in this context. Let us define
$C^{k}\coloneqq\{x\in\mathcal{H}\mid g(x)+\epsilon_{k}\leq 0\},$ (8)
[…]
Among them, we mention the algorithms in [fukushimaFinitelyConvergentAlgorithm1982, Iusem:1986a, iusemFinitelyConvergentIterative1987, DePierro:1988a], which were proposed for solving CFP. These methods employ the scheme in Eq.˜7 combined with operators related to $\bar{P}$ and $\widehat{P}$, where the projections onto the sets $C_{i}$, approximated from the inside by sets $C^{k}_{i}$, as in Eq.˜8, are replaced by projections onto hyperplanes separating $x^{k}$ from $C^{k}_{i}$. We stress here that the main Theorem in [fukushimaFinitelyConvergentAlgorithm1982, p. 1126], Theorem 1 in [iusemFinitelyConvergentIterative1987], and Lemma 1 in [DePierro:1988a] clearly state that only the tail of the sequence generated by these author’s algorithms are Fejér monotone, starting at some index which depends on the considered point. Therefore, the sequences are indeed Fejér* monotone.
@@
A similar situation occurs with the algorithm in [Behling:2024c], where the algorithm in [Iusem:1986a] is enhanced through the addition of a circumcentering step, using the machinery developed during the last few years in [Araujo:2022, Arefidamghani:2021, Arefidamghani:2023, Behling:2018, Behling:2018a, Behling:2020, Behling:2021a, Behling:2023, Behling:2024, Behling:2024b, Behling:2024c]. It is worthwhile to note that in the case of [Iusem:1986a, Behling:2024c, DePierro:1988a, iusemFinitelyConvergentIterative1987, fukushimaFinitelyConvergentAlgorithm1982], under the Slater hypothesis, if the $\epsilon_{k}$’s decrease to zero slowly enough, e.g., so that $\sum_{k=0}^{\infty}\epsilon_{k}=+\infty$ (take, for instance, $\epsilon_{k}=1/k$), convergence is indeed finite, i.e., there exists $\hat{k}$ such that $x^{\hat{k}}$ belongs to $C=\bigcap_{i=1}^{m}C_{i}$.
More recently, Kolobov:2021, Kolobov:2022 also make use of the perturbed approximation scheme Eq.˜7 in their algorithms to solve CFPs. In particular, Kolobov:2021 also shows that the tail of the sequence generated by their method is Fejér monotone. Thus, it is Fejér* monotone. Some other properties of Fejér* monotonicity related to optimization algorithms were explored recently by Arakcheev:2025, arakcheevFejerFejerMonotonicity2025, where Fejér* sequences were linked to Opial sequences, and in a survey on Fejér monotonicity by kohlenbachFejerMonotoneSequences2023. We refer the reader to these works for further details.
We highlight that the concept of Fejér* monotonicity is not novel, it has appeared especially when one is dealing with approximations of the feasible set from the inside, and in general is applied in the context of obtaning finite convergence results. Our work sheds some light on this concept, exploring its properties and differences with other weaker notions of Fejér monotonicity. This study will be carried out in the next sections.
@@ 3 Properties of Fejér* monotonicity
The first result shows some basic properties of Fejér* monotonicity. The result was already proved in [Behling:2024c, Prop. 2.5]. However, we chose to present the proof here, for the sake of completeness.
Proposition 3.1.
Let $(x^{k})_{k\in\mathds{N}}\subset\mathcal{H}$ be Fejér* monotone with respect to a nonempty set $M$ in $\mathcal{H}$. Then,
@@
    $(x^{k})_{k\in\mathds{N}}$ is Fejér* monotone with respect to $\operatorname{conv}(M)$.
Proof.
@@
Item (ii) is a direct consequence of Eq.˜2, since the $N(x)$-tail of sequence $(\left\lVert x^{k}-x\right\rVert)_{k\in\mathds{N}}$ is monotone and bounded, so the sequence converges.
For item (iii), take any $w\in\operatorname{conv}(M)$. Thus, $w$ can be written as $w=\sum_{i=1}^{p}\lambda_{i}w^{i}$, where $w^{i}\in M$, $\lambda_{i}\in[0,1]$ for all $i=1,2,\ldots,p$, $\sum_{i=1}^{p}\lambda_{i}=1$ and $p\in\mathds{N}$. Taking into account that $(x^{k})_{k\in\mathds{N}}$ is Fejér* monotone with respect to $M$, for each $i=1,2,\ldots,p$, there exists $N(w^{i})$ such that $\left\lVert x^{k+1}-w^{i}\right\rVert\leq\left\lVert x^{k}-w^{i}\right\rVert$ for all $k\geq N(w^{i})$. Squaring both sides of the last inequality and rearranging the terms, we get
$0$ $\leq\left\lVert{x^{k}}\right\rVert^{2}-\left\lVert{x^{k+1}}\right\rVert^{2}-2\left\langle x^{k}-x^{k+1},w^{i}\right\rangle.$ (9)
The above inequality also holds for all $k\geq K\coloneqq\max\{N(w^{1}),N(w^{2}),\ldots,N(w^{p})\}$. Multiplying Eq.˜9, for each $i=1,2,\ldots,p$, by the respective $\lambda_{i}$ and adding up, we obtain
$0$ $\leq\left\lVert{x^{k}}\right\rVert^{2}-\left\lVert{x^{k+1}}\right\rVert^{2}-2\left\langle x^{k}-x^{k+1},\sum_{i=1}^{p}\lambda_{i}w^{i}\right\rangle$ (10)
@@
and therefore, $\left\lVert x^{k+1}-w\right\rVert\leq\left\lVert x^{k}-w\right\rVert$, for all $k\geq K$. So $(x^{k})_{k\in\mathds{N}}$ is Fejér* monotone with respect to $\operatorname{conv}(M)$, and the proof is complete. ∎
Remark 3.2.
We remark that a Fejér monotone sequence with respect to a nonempty set $M\subset\mathcal{H}$ is also Fejér monotone with respect to $\operatorname{cl}\left({\operatorname{conv}(M)}\right)$, the closure of $\operatorname{conv}(M)$; see Cegielski:2012 and Combettes:2001a, or even Bauschke:2023. Furthermore, Proposition˜3.1(iii) guarantees that a Fejér* monotone sequence with respect to $M$ is as well Fejér* monotone with respect to $\operatorname{conv}(M)$. However, Fejér* monotonicity cannot always be extended to $\operatorname{cl}\left({\operatorname{conv}(M)}\right)$, as shown in the next example.
Example 3.3.
@@
$x^{2\ell+1}\coloneqq x^{2\ell}+(1/2^{\ell},0)\;,\quad x^{2\ell+2}\coloneqq\left(0,\sqrt{\|x^{2\ell+1}-(1,0)\|^{2}-1}\right).$ (12)
If $k$ is even, $x^{k+1}$ is obtained by moving $x^{k}$ horizontally to the right by an amount of $1/2^{(k/2)}$. If $k$ is odd, $x^{k+1}$ is the intersection between the vertical axis and the circle centered at $w^{0}=(1,0)$ passing through $x^{k}$. Figure˜1, illustrates the set $M$, the first nine points of sequence $(x^{k})_{k\in\mathds{N}}$, the convex hull $\operatorname{conv}(M)$, its closure $\operatorname{cl}\left({\operatorname{conv}(M)}\right)$ and the limit point $x^{*}$ of sequence $(x^{k})_{k\in\mathds{N}}$. The sequence that appears in Figure˜1 as well as all the calculations and the plot itself were coded and generated using Julia language [Bezanson:2017], on a MacbookAir M4, with 24GB RAM and 10-core CPU.
Figure 1: From Example˜3.3: set $M$, first nine points of sequence $(x^{k})_{k\in\mathds{N}}$, $\operatorname{conv}(M)=(0,1]\times\{0\}$, $\operatorname{cl}\left({\operatorname{conv}(M)}\right)\setminus\operatorname{conv}(M)=\{(0,0)\}$ and limit point $x^{*}$ of sequence $(x^{k})_{k\in\mathds{N}}$.
We now show that $(x^{k})_{k\in\mathds{N}}$ is Fejér* monotone with respect to $M$. Denote $x^{k}=(\alpha_{k},\beta_{k})$, $k\in\mathds{N}$. One can see that $1\leq\beta_{k+1}\leq\beta_{k}$ as well as $\|x^{k}-w^{0}\|\geq 1$, for all $k\in\mathds{N}$, so that the sequence $(x^{k})_{k\in\mathds{N}}$ is well-defined.